# AutoLibra: Agent Metric Induction from Open-Ended Feedback

### v1

@@ Acknowledgment
This work is supported by ONR grant N000142412532 and NSF grant IIS-2247357. We thank Google Cloud Platform and Modal Platform for their credits. We thank Yutong Zhang, Hayley Zhang, Yijia Shao, Michelle S Lam, Manling Li, Ryan Louie, Yanzhe Zhang, Xuhui Zhou, Maarten Sap, Sherry Tongshuang Wu, Shikhar Murty, Saujas Vaduguru, Chenghao Yang, Xizhi Xiao, Anant Sinha and all members of Stanford SALT Lab for their help and feedback throughout this project.
@@
    • A Algorithm of AutoLibra Ladder
    • B AutoLibra Ladder Experiment Configuration
@@
    • C Rules and Environment Details
    • D Experiment Results
    • E Metric Scores
    • F Metric Examples
    • G Prompts
    • H Qualitative Observations of Agent Performance
@@
    • I Rules and Environment Details
    • J Experiment Results
    • K Metric Scores
    • L Metric Examples
    • M Prompts
    • N Qualitative Observations of Agent Performance
@@
    • O NNetNav-Live Induced Metrics
Appendix A Algorithm of AutoLibra Ladder Experiment
@@
1:for $i$ in range($n\_iters$do
2:     for $task$ in $selected\_tasks$ do
@@
6:     $metrics_{i}=AutoLibra.extract_{m}etrics(traj_{i},annotations_{i})$
7:     $traj\_scores_{i}=AutoLibra.llm_{e}val(metrics_{i},traj_{i})$
@@
11:         for $task$ in $selected\_tasks$ do
12:              _$\_,curr_{s}cores=agent.play(task,prompt)$
@@
Appendix B AutoLibra Ladder Experiment Setup
@@
Appendix C Baba-is-ai Rules and Environment Configuration
@@
Appendix D Baba-is-ai Experiment Results
@@
Appendix E Baba-is-ai Metric Scores
@@
Appendix F Baba-is-ai Metric Examples
@@
Appendix G Baba-is-ai Prompts
[…]
[ backgroundcolor=lightgreen, linecolor=darkgreen, linewidth=2pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, leftmargin=0pt, rightmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt, splittopskip=litbottomskip=10pt, frametitle=Iteration 0 Baba-is-ai Prompt, frametitlebackgroundcolor=darkgreen, frametitlefont=, frametitlerule=true, frametitleaboveskip=8pt, frametitlebelowskip=8pt, ]GreenBox
@@ Appendix G Baba-is-ai Prompts
rule ball 8 steps to the right and 2 steps up
rule is 9 steps to the right and 2 steps up
rule win 10 steps to the right and 2 steps up
rule wall 1 step up
rule is 1 step to the right and 1 step up
rule stop 2 steps to the right and 1 step up
wall 5 steps to the right and 1 step up
rule door 10 steps to the right and 1 step up
wall 5 steps to the right
ball 6 steps to the right
wall 5 steps to the right and 1 step down
wall 5 steps to the right and 2 steps down
wall 5 steps to the right and 3 steps down
rule baba 4 steps down
rule is 1 step to the right and 4 steps down
rule you 2 steps to the right and 4 steps down
wall 5 steps to the right and 4 steps down
door 7 steps to the right and 4 steps down
@@
rule is 7 steps to the right and 3 step up
rule win 8 steps to the right and 3 step up
rule wall 2 step to the left and 2 step up
rule is 1 step to the left and 2 step up
rule stop 2 step up
wall 3 steps to the right and 2 step up
wall 3 steps to the right and 1 step up
key 4 steps to the right and 1 step up
wall 3 steps to the right
rule ball 7 steps to the right
wall 3 steps to the right and 1 step down
wall 3 steps to the right and 2 steps down
rule baba 2 step to the left and 3 steps down
rule is 1 step to the left and 3 steps down
rule you 3 steps down
wall 3 steps to the right and 3 steps down
ball 4 steps to the right and 3 steps down
@@
wall is stop
baba is you
Objects on the map:
rule wall at (1, 1)
rule is at (2, 1)
rule stop at (3, 1)
rule is at (7, 3)
rule win at (8, 3)
rule ball at (7, 0)
rule baba at (5, 6)
rule is at (6, 6)
rule you at (7, 6)
wall at (4, 0), (4, 1), (4, 2), (4, 3), (4, 4), (4, 5), (4, 6)
key at (5, 1)
ball at (5, 6)
@@
Active rules:
baba is you
@@
Appendix H Qualitative Observations of baba-is-ai Agent Performance
The metrics induced by AutoLibra (Table 4) capture the behavior of the agent effectively, with coverage increasing from 65% at Iteration 0 to 92% at Iteration 3 and a low mean redundancy of 56%. Notably, earlier metrics were found to describe higher-level behaviors than later metrics, reflective of the more complex behaviors that the agent demonstrated in later iterations, as well as the compositionality of induced metrics. Code changes were selected to specifically target given metrics, and as seen in Figure 8, the agent’s performance on the targeted metric improved significantly in the iteration following the code change. This demonstrates the utility of AutoLibra for fine-grained agent improvement, as well as the human interpretability of the induced metrics.
@@ Appendix H Qualitative Observations of baba-is-ai Agent Performance
  • Augmentation of single-shot example with movement instructions, targeting improvement of Rule Manipulation and Execution [Uncaptioned image]
@@
Appendix I MiniHack Rules and Environment Configuration
@@
Similar to our experiments with baba-is-ai, the Ladder improvement process for MiniHack also follows the algorithm detailed in Appendix A. Two full iterations of agent improvement with AutoLibra were performed on MiniHack. Four representative tasks for MiniHack are used in iterative metric improvement, with the remainder held out for evaluation. The agent is evaluated on the MiniHack environment and any changes to the agent code at the beginning of each iteration, and the environment score, trajectory performance, and other metrics are recorded at the end of each iteration. GPT-4o-241120 is used as the agent model.
@@
Appendix J MiniHack Experiment Results
@@
Appendix K MiniHack Metric Performance
Iteration 0 1 2
Target Navigation Effectiveness [Uncaptioned image] 16.67% 8.33% 41.67%
Efficient Exploration and Map Memory Utilization [Uncaptioned image] 16.67% 0.00% 25.00%
Hazard Awareness and Equipment Utilization [Uncaptioned image] 0.00% 0.00% 0.00%
Boulder Manipulation Strategy [Uncaptioned image] 0.00% 0.00% 0.00%
Combat Engagement and Survival [Uncaptioned image] 8.33% 8.33% 25.00%
Role-Specific Ability Utilization [Uncaptioned image] 0.00% 0.00% 0.00%
Spatial Awareness and Interpretation [Uncaptioned image] - 16.67% 58.33%
Object Pickup Efficiency [Uncaptioned image] - 0.00% 0.00%
Giant Rats Encounter Handling [Uncaptioned image] - - 25.00%
         Coverage 82.89% 81.82% 87.84%
Redundancy 61.11% 65.63% 71.30%
Table 8: Metric Performance for MiniHack AutoLibra Iterations 0–2
Appendix L MiniHack Metric Examples
Iteration 0: Target Navigation Effectiveness Explanation: This metric evaluates the agent’s ability to accurately identify and navigate toward key goal locations (e.g., the downward stairs). Positive behaviors show clear planning and directional focus, while negative examples reveal distractions or failures in goal-oriented movement. Good Behaviors: Behavior: The agent decides to move south and reaches the downward stairs. Behavior: The agent’s reasoning shows it plans its action to reach the stairs. Behavior: From its reasoning, the agent appears to recognize the goal of locating the stairs downward. Behavior: From the trajectory and reasoning, the agent consistently aimed toward the general direction of its goal—the stairs down. Bad Behaviors: Behavior: The repeated attempts to kick the goblin instead of progressing towards the stairs. Behavior: The agent repeatedly interacts with adjacent stairs up but does not visibly attempt to identify or locate downward stairs. Behavior: The agent explored the map but did not identify the downward stair in its trajectory. Behavior: Throughout the trajectory, the agent did not manage to locate the downward stairs, despite exploring various places in the environment. Behavior: The agent spends multiple actions traversing the map without making meaningful progress toward the downward stairs. Behavior: Performed in the sequence of actions where the agent explores but eventually doesn’t find the stairway downward, ending the session unsuccessfully. Behavior: During its navigation, the agent didn’t succeed in reaching the stairs down.
Appendix M MiniHack Prompts
[…]
[ backgroundcolor=lightgreen, linecolor=darkgreen, linewidth=2pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, leftmargin=0pt, rightmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt, splittopskip=litbottomskip=10pt, frametitle=Iteration 0 MiniHack Prompt, frametitlebackgroundcolor=darkgreen, frametitlefont=, frametitlerule=true, frametitleaboveskip=8pt, frametitlebelowskip=8pt, ]MyGreenBox
{MyGreenBox}
[frametitle=Iteration 0 MiniHack Prompt] You are an agent playing MiniHack. The following are the possible actions you can take in the game, followed by a short description of each action:
  • north: move north,
  • east: move east,
  • south: move south,
  • west: move west,
  • northeast: move northeast,
  • southeast: move southeast,
  • southwest: move southwest,
  • northwest: move northwest,
  • far north: move far north,
  • far east: move far east,
  • far south: move far south,
  • far west: move far west,
  • far northeast: move far northeast,
  • far southeast: move far southeast,
  • far southwest: move far southwest,
  • far northwest: move far northwest,
  • up: go up the stairs,
  • down: go down the stairs,
  • wait: rest one move while doing nothing,
  • more: display more of the message,
  • apply: apply (use) a tool,
  • close: close an adjacent door,
  • open: open an adjacent door,
  • eat: eat something,
  • force: force a lock,
  • kick: kick an enemy or a locked door or chest,
  • loot: loot a box on the floor,
  • pickup: pick up things at the current location if there are any,
  • pray: pray to the gods for help,
  • puton: put on an accessory,
  • quaff: quaff (drink) something,
  • search: search for hidden doors and passages,
  • zap: zap a wand
For task which name consists "corridor": Your goal is to explore the level and reach the stairs down.
For task which name consists “quest”: Your goal is to explore the level, fight monsters, and navigate rooms and mazes to ultimately reach the stairs down.
For task which name consists “boxoban”: You are playing Boxoban, a box-pushing game inspired by Sokoban. Your goal is to push the boulders onto the fountains on the map. You can push the boulders by walking into them, as long as there are no obstacles behind them.
For task which name consists “mazewalk”: Your goal is to explore the level and reach the stairs down.
Otherwise: Your goal is to get as far as possible in the game.
Additional Feedback (if provided): I will also present lists of positive and negative feedback based on your previous attempts on this task, use this information to change your strategy and improve your performance.  Positive: positive Negative: negative
PLAY!
{MyGreenBox}
[frametitle=Iteration 1 MiniHack Prompt] You are an agent playing MiniHack. The following are the possible actions you can take in the game, followed by a short description of each action:
  • north: move north,
  • east: move east,
  • south: move south,
  • west: move west,
  • northeast: move northeast,
  • southeast: move southeast,
  • southwest: move southwest,
  • northwest: move northwest,
  • far north: move far north,
  • far east: move far east,
  • far south: move far south,
  • far west: move far west,
  • far northeast: move far northeast,
  • far southeast: move far southeast,
  • far southwest: move far southwest,
  • far northwest: move far northwest,
  • up: go up the stairs,
  • down: go down the stairs,
  • wait: rest one move while doing nothing,
  • more: display more of the message,
  • apply: apply (use) a tool,
  • close: close an adjacent door,
  • open: open an adjacent door,
  • eat: eat something,
  • force: force a lock,
  • kick: kick an enemy or a locked door or chest,
  • loot: loot a box on the floor,
  • pickup: pick up things at the current location if there are any,
  • pray: pray to the gods for help,
  • puton: put on an accessory,
  • quaff: quaff (drink) something,
  • search: search for hidden doors and passages,
  • zap: zap a wand
For task name contains “corridor”:
@@
For task name contains “quest”:
Your goal is to explore the level, fight monsters, and navigate rooms and mazes to ultimately reach the stairs down.
@@
For task name contains “boxoban”:
You are playing Boxoban, a box-pushing game inspired by Sokoban. Your goal is to push the boulders onto the fountains on the map. You can push the boulders by walking into them, as long as there are no obstacles behind them.
@@
For task name contains “mazewalk”:
Your goal is to explore the level and reach the stairs down.
@@
Otherwise:
Your goal is to get as far as possible in the game.
@@
For task name contains “corridor”:
Your goal: Your goal is to explore the level and reach the stairs down.
@@
For task name contains “quest”:
Your goal: Your goal is to explore the level, fight monsters, and navigate rooms and mazes to ultimately reach the stairs down.
@@
For task name contains “boxoban”:
Your goal: You are playing Boxoban, a box-pushing game inspired by Sokoban. Your goal is to push the boulders onto the fountains on the map. You can push the boulders by walking into them, as long as there are no obstacles behind them.
@@
For task name contains “mazewalk”:
Your goal: Your goal is to explore the level and reach the stairs down.
@@
Absolute Wall-Following Protocol:
Core Principle: “Follow priority order, probe dark directions ONLY when they’re current priority.”
Movement Validation:
– Priority rotation ONLY occurs AFTER SUCCESSFUL MOVEMENT.
– Wall collisions preserve current priority order.
@@
Walkthrough Example:
Initial: [E,N,W,S]
1. Move E until wall. 2. At start: E (wall) → N (open) → move north (no rotation). 3. Next: E (open) → move east (rotate to SENW).
Otherwise:
Your goal: Your goal is to get as far as possible in the game.
@@
Appendix N Qualitative Observations of MiniHack Agent Performance
@@
The induced metrics and the agent’s per-task performance are shown in Table 9 and Table 8, respectively. A substantial improvement in the agent’s task completion performance is observed from Iteration 1 to Iteration 2, with the agent achieving both a higher environment score and trajectory performance on the held-out tasks compared to the baseline agent. Specifically, the agent’s performance on metrics increased correspondingly to code changes, like [Uncaptioned image] improving from $16.67\%$ to $41.67\%$, demonstrating the utility of AutoLibra for fine-grained agent improvement. This is discussed in more detail in the following sections.
N.1 Extracted Metrics and Improvements
The metrics induced by AutoLibra (Table 9) capture the behavior of the agent through all iterations, with coverage of 83% at Iteration 0 to 88% at Iteration 2. Due to the diversity of tasks, metrics pertain to one or two tasks, matching the expectation that AutoLibra should generate fine-grained metrics. Notably, the metrics generated at Iteration 0 are nearly comprehensive as they demonstrate good coverage of all four MiniHack tasks, while the later metrics were found to specifically describe detailed behaviors for targeting one task each. Code changes were selected to specifically target improvements in given metrics, and as seen in Table 8, the agent’s performance on the targeted metric improved significantly in the iteration following the code change. This demonstrates the utility of AutoLibra for fine-grained agent improvement, as well as the human interpretability of the induced metrics. On the other hand, some metrics remain unchanged over iterations. The unchanged metrics are [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. These metrics are all tightly related to Boxoban or Quest environments. As these two environments contain high level of randomness, interactions with system, and sophisticated planning, our prompt and example based guidance do not provide obvious improvements. This result also indicates that simply coding might not be sufficient if the gap between the agent’s capability and the task’s complexity is too significant.
@@
N.2 Held-Out Task Performance
@@
Appendix O NNetNav-Live Induced Metrics

### v3

@@ Acknowledgments
This work is supported by ONR grant N000142412532, and NSF grant IIS-2247357, and DARPA grant Friction for Accountability in Conversational Transactions. We thank Google Cloud Platform and Modal Platform for their credits. We thank Yutong Zhang, Hayley Zhang, Yijia Shao, Michelle S Lam, Manling Li, Ryan Louie, Yanzhe Zhang, Xuhui Zhou, Maarten Sap, Sherry Tongshuang Wu, Shikhar Murty, Saujas Vaduguru, Chenghao Yang, Xizhi Xiao, Anant Sinha and all members of Stanford SALT Lab for their help and feedback throughout this project.
@@
    • A Illustration of AutoLibra Method
    • B Algorithm of AutoLibra Ladder
    • C AutoLibra Ladder Experiment Configuration
@@
    • D Rules and Environment Details
    • E Experiment Results
    • F Metric Scores
    • G Metric Examples
    • H Prompts
    • I Qualitative Observations of Agent Performance
@@
    • J Rules and Environment Details
    • K Experiment Results
    • L Metric Scores
    • LABEL:appendix:minihack_metrics Metric Examples
    • LABEL:appendix:minihack_prompts Prompts
    • O Qualitative Observations of Agent Performance
@@
    • P NNetNav-Live Induced Metrics
Appendix A Illustration of AutoLibra Method
Figure 6: Feedback Grounding
Figure 7: Behavior Clustering
Figure 8: Evaluation with induced metrics
Figure 9: Meta evaluation
Appendix B Algorithm of AutoLibra Ladder Experiment
@@
1:for $i$ in range(n_itersn\_iters) do
2:  for tasktask in selected_tasksselected\_tasks do
@@
6:  $metrics_{i}=AutoLibra.extract\_metrics(traj_{i},annotations_{i})$
7:  $traj\_scores_{i}=AutoLibra.llm\_eval(metrics_{i},traj_{i})$
@@
11:   for tasktask in selected_tasksselected\_tasks do
12:     _,curr_scores=agent.play(task,prompt)\_,curr\_scores=agent.play(task,prompt)
@@
Appendix C AutoLibra Ladder Experiment Setup
@@
Appendix D Baba-is-ai Rules and Environment Configuration
@@
Appendix E Baba-is-ai Experiment Results
@@
Appendix F Baba-is-ai Metric Scores
@@
Appendix G Baba-is-ai Metric Examples
@@
Appendix H Baba-is-ai Prompts
@@
rule ball 8 steps to the right and 2 steps up rule is 9 steps to the right and 2 steps up rule win 10 steps to the right and 2 steps up rule wall 1 step up rule is 1 step to the right and 1 step up rule stop 2 steps to the right and 1 step up wall 5 steps to the right and 1 step up rule door 10 steps to the right and 1 step up wall 5 steps to the right ball 6 steps to the right wall 5 steps to the right and 1 step down wall 5 steps to the right and 2 steps down wall 5 steps to the right and 3 steps down rule baba 4 steps down rule is 1 step to the right and 4 steps down rule you 2 steps to the right and 4 steps down wall 5 steps to the right and 4 steps down door 7 steps to the right and 4 steps down
@@
rule is 7 steps to the right and 3 step up rule win 8 steps to the right and 3 step up rule wall 2 step to the left and 2 step up rule is 1 step to the left and 2 step up rule stop 2 step up wall 3 steps to the right and 2 step up wall 3 steps to the right and 1 step up key 4 steps to the right and 1 step up wall 3 steps to the right rule ball 7 steps to the right wall 3 steps to the right and 1 step down wall 3 steps to the right and 2 steps down rule baba 2 step to the left and 3 steps down rule is 1 step to the left and 3 steps down rule you 3 steps down wall 3 steps to the right and 3 steps down ball 4 steps to the right and 3 steps down
@@
wall is stop baba is you
Objects on the map: rule wall at (1, 1) rule is at (2, 1) rule stop at (3, 1) rule is at (7, 3) rule win at (8, 3) rule ball at (7, 0) rule baba at (5, 6) rule is at (6, 6) rule you at (7, 6) wall at (4, 0), (4, 1), (4, 2), (4, 3), (4, 4), (4, 5), (4, 6) key at (5, 1) ball at (5, 6)
@@
Active rules: baba is you
@@
Appendix I Qualitative Observations of baba-is-ai Agent Performance
The metrics induced by AutoLibra capture the behavior of the agent effectively, with coverage increasing from 65% at Iteration 0 to 92% at Iteration 3 and a low mean redundancy of 56%. Notably, earlier metrics were found to describe higher-level behaviors than later metrics, reflective of the more complex behaviors that the agent demonstrated in later iterations, as well as the compositionality of induced metrics. Code changes were selected to specifically target given metrics, and the agent’s performance on the targeted metric improved significantly in the iteration following the code change. This demonstrates the utility of AutoLibra for fine-grained agent improvement, as well as the human interpretability of the induced metrics.
@@ Appendix I Qualitative Observations of baba-is-ai Agent Performance
  • Augmentation of single-shot example with movement instructions, targeting improvement of
Execution [Uncaptioned image]
@@
Appendix J MiniHack Rules and Environment Configuration
@@
Similar to our experiments with baba-is-ai, the Ladder improvement process for MiniHack also follows the algorithm detailed in Appendix B. Two full iterations of agent improvement with AutoLibra were performed on MiniHack. Four representative tasks for MiniHack are used in iterative metric improvement, with the remainder held out for evaluation. The agent is evaluated on the MiniHack environment and any changes to the agent code at the beginning of each iteration, and the environment score, trajectory performance, and other metrics are recorded at the end of each iteration. GPT-4o-241120 is used as the agent model.
@@
Appendix K MiniHack Experiment Results
@@
Appendix L MiniHack Metric Performance
@@
For task name contains “quest”: Your goal is to explore the level, fight monsters, and navigate rooms and mazes to ultimately reach the stairs down.
@@
For task name contains “boxoban”: You are playing Boxoban, a box-pushing game inspired by Sokoban. Your goal is to push the boulders onto the fountains on the map. You can push the boulders by walking into them, as long as there are no obstacles behind them.
@@
For task name contains “mazewalk”: Your goal is to explore the level and reach the stairs down.
@@
Otherwise: Your goal is to get as far as possible in the game.
@@
For task name contains “corridor”: Your goal: Your goal is to explore the level and reach the stairs down.
@@
For task name contains “quest”: Your goal: Your goal is to explore the level, fight monsters, and navigate rooms and mazes to ultimately reach the stairs down.
@@
For task name contains “boxoban”: Your goal: You are playing Boxoban, a box-pushing game inspired by Sokoban. Your goal is to push the boulders onto the fountains on the map. You can push the boulders by walking into them, as long as there are no obstacles behind them.
@@
For task name contains “mazewalk”: Your goal: Your goal is to explore the level and reach the stairs down.
@@
Absolute Wall-Following Protocol: Core Principle: “Follow priority order, probe dark directions ONLY when they’re current priority.” Movement Validation: – Priority rotation ONLY occurs AFTER SUCCESSFUL MOVEMENT. – Wall collisions preserve current priority order.
@@
Walkthrough Example: Initial: [E,N,W,S] 1. Move E until wall. 2. At start: E (wall) → N (open) → move north (no rotation). 3. Next: E (open) → move east (rotate to SENW).
Otherwise: Your goal: Your goal is to get as far as possible in the game.
@@
Appendix O Qualitative Observations of MiniHack Agent Performance
@@
The induced metrics and the agent’s per-task performance are shown in Table 7 and Table LABEL:tab:metric_mini_perf, respectively. A substantial improvement in the agent’s task completion performance is observed from Iteration 1 to Iteration 2, with the agent achieving both a higher environment score and trajectory performance on the held-out tasks compared to the baseline agent. Specifically, the agent’s performance on metrics increased correspondingly to code changes, like [Uncaptioned image] improving from $16.67\%$ to $41.67\%$, demonstrating the utility of AutoLibra for fine-grained agent improvement. This is discussed in more detail in the following sections.
O.1 Extracted Metrics and Improvements
The metrics induced by AutoLibra (Table 7) capture the behavior of the agent through all iterations, with coverage of 83% at Iteration 0 to 88% at Iteration 2. Due to the diversity of tasks, metrics pertain to one or two tasks, matching the expectation that AutoLibra should generate fine-grained metrics. Notably, the metrics generated at Iteration 0 are nearly comprehensive as they demonstrate good coverage of all four MiniHack tasks, while the later metrics were found to specifically describe detailed behaviors for targeting one task each. Code changes were selected to specifically target improvements in given metrics, and as seen in Table LABEL:tab:metric_mini_perf, the agent’s performance on the targeted metric improved significantly in the iteration following the code change. This demonstrates the utility of AutoLibra for fine-grained agent improvement, as well as the human interpretability of the induced metrics. On the other hand, some metrics remain unchanged over iterations. The unchanged metrics are [Uncaptioned image] , [Uncaptioned image] , [Uncaptioned image] , and [Uncaptioned image] . These metrics are all tightly related to Boxoban or Quest environments. As these two environments contain high level of randomness, interactions with system, and sophisticated planning, our prompt and example based guidance do not provide obvious improvements. This result also indicates that simply coding might not be sufficient if the gap between the agent’s capability and the task’s complexity is too significant.
@@
O.2 Held-Out Task Performance
@@
Appendix P NNetNav-Live Induced Metrics